A rocket consists of a right circular cylinder of height 20 m surmounted by a cone whose height and diameter are equal and whose radius is the same as that of the cylindrical section. What should this radius be (rounded to two decimal places) if the total volume is to be
700𝜋/3 m^3?

Answer by mananth(16946)


.
Volume of cone +volume of cylinder = Total volume
2/3 pi r^3 + 40pir^2= 700 pi/3
2/3 r^3+40r^2 = 700/3
2r^3+120r^2 = 700
r^3+60r^2-700=0
Use calculator to solve
r=2.369